OpenStudy (anonymous):

prove the identity. do not skip any steps. every step must flow directly from the step before. sinx=cosx/cotx

jimthompson5910 (jim_thompson5910):

cot(x) = cos(x)/sin(x)

OpenStudy (anonymous):

that's all?

jimthompson5910 (jim_thompson5910):

it's a hint to get you started

jimthompson5910 (jim_thompson5910):

let me know how far you get

OpenStudy (anonymous):

ok.. one sec

OpenStudy (jdoe0001):

"prove the identity. do not skip any steps. every step must flow directly from the step before." <---- sounds like instructions to the paper holder

OpenStudy (jdoe0001):

anyhow as jim_thompson5910 said, use that, simplify, be happy, eat ice-cream

OpenStudy (anonymous):

next step is cotx=cosx^2 sinx?

jimthompson5910 (jim_thompson5910):

you have this so far, right? \[\Large \sin(x) = \frac{\cos(x)}{\cot(x)}\] \[\Large \sin(x) = \frac{\cos(x)}{\cos(x)/\sin(x)}\]

OpenStudy (anonymous):

no i didn't.. i am really struggling with this.. but i'm really trying to understand.. im thinking the next step is the cosx cancel out and leaves sinx=sinx

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (jdoe0001):

hint: \(\bf \cfrac{\frac{a}{b}}{\frac{c}{{\color{blue}{ d}}}}\implies \cfrac{a}{b}\cdot \cfrac{{\color{blue}{ d}}}{c}\)

jimthompson5910 (jim_thompson5910):

this is what the full step by step picture looks like \[\Large \sin(x) = \frac{\cos(x)}{\cot(x)}\] \[\Large \sin(x) = \frac{\cos(x)}{\cos(x)/\sin(x)}\] \[\Large \sin(x) = \cos(x)*\frac{\sin(x)}{\cos(x)}\] \[\Large \sin(x) = \cancel{\cos(x)}*\frac{\sin(x)}{\cancel{\cos(x)}}\] \[\Large \sin(x) = \sin(x)\]

OpenStudy (anonymous):

but sinx=sinx does not equal the original problem? I thought the question was wanting me to show how they are the same

OpenStudy (anonymous):

i think i was wrong. it says probe and i guess thats what you did. is there more to the problem?

jimthompson5910 (jim_thompson5910):

once you get to sin(x) = sin(x), you have shown that the two original sides are the same expression

OpenStudy (anonymous):

alright thank you for your help. would you mind checking two problems that I already did?

jimthompson5910 (jim_thompson5910):

that effectively proves/verifies the identity

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

(sinx+cosx)^2=+sinx cosx [sinx+cosx]^2 =sin^2+2sinx cosx+cos^2x =[sin^2x+cos^2x]+2sinx cosx sin^2x+cos^2x=1 by the pythagorean identity [sin^2x+cos^2x]+2sinx cosx=1+2sinx cosx

OpenStudy (anonymous):

@jim_thompson5910 are you stilll here?

jimthompson5910 (jim_thompson5910):

is there a typo in your first line when you wrote "(sinx+cosx)^2=+sinx cosx" ?

OpenStudy (anonymous):

oh yes i'm sorry.. my num lock was on.. it's (sinx+cosx)^2=1+2sinx cosx

jimthompson5910 (jim_thompson5910):

it looks good, I would do it like this (see attached)

jimthompson5910 (jim_thompson5910):

but you effectively have the right idea

OpenStudy (anonymous):

thank you.. total i have 3 more problems.. I was able to do one, could you check that and help with the other two? ill attempt them myself again though

OpenStudy (anonymous):

the one i did was 1+cos(8x)=2cos^2(4x) cos(A+B)=cosAcosB-sinAsinB A=B=4x cos8x=cos^2 4x-sin^2 4x =cos^2 4x-(-1-cos^2 4x)=2cos^2 4x-1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

looks good, the +1 in the left side of the original equation will cancel with the -1 in 2cos^2 (4x)-1

OpenStudy (anonymous):

so it would just be 2cos^2 4x

OpenStudy (anonymous):

now the two problems i am having trouble with.. sinx(cosx cotx-sinx)=cos(2x)

jimthompson5910 (jim_thompson5910):

recall that cot(x) = cos(x)/sin(x)

jimthompson5910 (jim_thompson5910):

distribute the sine through. What do you get?

OpenStudy (anonymous):

cotx(sinx)=cosx?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

the sines will cancel

OpenStudy (anonymous):

then i have cotx=cosx

jimthompson5910 (jim_thompson5910):

cotx=cosx is not an identity

jimthompson5910 (jim_thompson5910):

cotx(sinx)=cosx is an identity

OpenStudy (anonymous):

so that's the end of it? all i have is cotx=cosx/sinx cotx(sinx)=cosx?

jimthompson5910 (jim_thompson5910):

that's just part of it to help you do the whole thing

OpenStudy (anonymous):

ok so after cotx(sinx)=cosx i will have cotx=cosx

jimthompson5910 (jim_thompson5910):

sinx(cosx cotx-sinx)=cos(2x) sinx(cosx cotx) + sinx*(-sinx)=cos(2x) cosx*cosx - sinx*sinx=cos(2x) what's next?

OpenStudy (anonymous):

wouldn't it be cosx^2 sinx^2=cos(2x)

jimthompson5910 (jim_thompson5910):

i think you mean cos^2(x) - sin^2(x) = cos(2x)

OpenStudy (anonymous):

im sorry. im really trying to understand this.. is that the end of that problem?

jimthompson5910 (jim_thompson5910):

no, the ultimate end is where the two sides are the same expression

OpenStudy (anonymous):

OH IT FINALLY CLICKED! i understan now,

jimthompson5910 (jim_thompson5910):

here is a nontrig example of an identity (and proving/verifying it) 2x+3x = 5x 5x = 5x

OpenStudy (anonymous):

soooo cos^2x-sin^2x=cos(2x) then we need to get -sin^2x away... can i just add +sin^2x to leav with cos^2x=cos(2x)

jimthompson5910 (jim_thompson5910):

no, you cannot add things to both sides you must only manipulate one side only

jimthompson5910 (jim_thompson5910):

and cos^2x=cos(2x) is false

OpenStudy (anonymous):

cos^2x-sin^2x=cos(2x) cotx=cos^2x/-sin^2x=cos(2x) which would then give me cos(2x) on both sides

OpenStudy (anonymous):

i feel like im messing up more now @jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910 please help me finish this problem

jimthompson5910 (jim_thompson5910):

one sec and I'll type something up

jimthompson5910 (jim_thompson5910):

See attached

OpenStudy (anonymous):

my very last problem is different.. show all of the work to derive a real number value that will make this a true statement: if tanx=-5/12 and x is in quadrant 2, then sin 2x=______

OpenStudy (anonymous):

i would plug -5/12 in for x.... to be.. sin 2(-5/12) then solve right

OpenStudy (anonymous):

@jim_thompson5910 sin(-5/12)=-.404

jimthompson5910 (jim_thompson5910):

no that's not correct

OpenStudy (anonymous):

did i mess up from the beginning or just the -.404

jimthompson5910 (jim_thompson5910):

if tan(x) = -5/12, then what is x?

OpenStudy (anonymous):

isnt it -5/12

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):